The orbital time of Neptune around the Sun is 165 Earth years and its orbital path is 30 times longer than that of the Earth.


a Calculate the orbital speed of Neptune relative to the Earth. orbital speed b


State how the speed of travel of Neptune compares with that of the Earth.


1
Expert's answer
2022-05-31T16:11:34-0400

"v_2=\frac{30\cdot 2\pi a}{T_2}=1.14~\frac{au}y,"

"v_1=\frac{2\pi a}{T_1}=6.28~\frac {au}y,"

"\frac{v_1}{v_2}=5.5."